ELECTRONIC SPECTRA AND IONIZATION POTENTIALS FOR UO AND UO_{2}

\underline{VASILIY V. GONCHAROV}, LEONID A. KALEDIN, JIANDE HAN AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The electronic spectra and ionization potentials of UO and UO_{2} are being investigated using resonantly enhanced multiphoton ionization combined with mass resolved ion detection. Several new band systems for both molecules have been observed in the near UV spectral range (330-370 nm). The ionization potentials have been probed using single- and two-photon (1+1’) excitation. The spectroscopic data for UO_{2} have been used to evaluate the predictions of ab initio and ligand field theory models. A summary of the spectroscopic results and a discussion of the electronic state assignments will be presented.